OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school?

OpenStudy (anonymous):

g girls b boys g+b=89 b=2g-8

OpenStudy (anonymous):

g girls b boys g+b=89 b=2g+8 @protosbg isnt it supossed to a + ?

OpenStudy (anonymous):

boys are more

OpenStudy (anonymous):

the number of boys is 8 more (+8), twice (2g) the number of girls. Is my logic wrong here?

OpenStudy (anonymous):

I'm so confused right now lol

OpenStudy (anonymous):

:D yes yes

OpenStudy (anonymous):

sorry :)

OpenStudy (anonymous):

lol relax Bianca will solve it.

OpenStudy (anonymous):

so it a +. Do you agree @protosbg ?

OpenStudy (anonymous):

:D absolutely 10x for correction

OpenStudy (anonymous):

lol :D sweet. Do you get it @biancataylor515 ?

OpenStudy (anonymous):

I understand what you did, but i'm not so sure on how to solve it..

OpenStudy (anonymous):

substitute g from first equation in second :)

OpenStudy (anonymous):

yep just do protosbg says and tell me what answer you get.

OpenStudy (anonymous):

62 @biancataylor515 ?

OpenStudy (anonymous):

i dont know i stop here :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

eqn 1 g= 89-b eqn 2 b=2g+8 b=2(89-b) + 8 substitute eqn 1 into eqn 2 b=178-2b+8 b=186-2b 3b=186 b=62

OpenStudy (anonymous):

okay i got the same thing a while ago sorry i didn't reply sooner but thank you so so so so much :)